[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Aug 24, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed Aug 24 13:55:37 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
The TCMWindTool2016 was tested and working fine now.

The AWIPS lp1 printer working again. It was powered off and on and now,  
which fixed the problem. Forecasters are reminded that they can select lp2  
to print from AWIPS in the future if there are issues with lp1."
HSU:
Remember to continue to follow Mike's guidelines on ATCF cleanup.

Mark is coordinating with CIRA on the differences between the TCGI run at  
CIRA and on WCOSS.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160824/77ec8b6e/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list